1. 12-4n=4
2. 1/3(s=6)=3
3. 1/5(t-2)=0
4. -5t - 5= -5

Answer by checkley71(8403)

12-4N=4 OR -4N=-8 OR N=2
1/3(S-6)=3 OR S/3-2=3 OR S/3=5 OR S=15
1/5(T-2)=0 OR T/5-2/5=0 OR (T-2)/5=0 OR T-2=0 OR T=2
-5T-5=-5 OR -5T=0 OR T=0